OpenStudy (anonymous):

Give the slope of each line & then determine whether the 2 lines are parallel, perpendicular or neither. 3x+4y=8 & 12x+16y=0

OpenStudy (anonymous):

put them both in the form y=mx+b, if m is equal for both, they are parallel. If you can flip m of one equation and change it's sign and it equals the m of the other, they are perpendicular. If neither of these conditions exists, it's neither parallel or perpendicular

OpenStudy (anonymous):

I am terrible at these that is why I am asking for help. I just seem to not be able to work these very good.

OpenStudy (anonymous):

this locked up on me again.

OpenStudy (anonymous):

For perpendicular check, you multiply both m's, if they equal -1 they are perpendicular, do u know how to put them into y=mx+b format?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

move y to one side and isolate it....3x+4y=8... 4y=8-3x...y=(8-3x)/4...y=2-3/4x...y=-3/4x+2....the number atached to x is your m

OpenStudy (anonymous):

ok still lost

OpenStudy (anonymous):

is that a fraction or is it -3 divided by 4x?

OpenStudy (anonymous):

it's (-3/4)x

OpenStudy (anonymous):

ok so (-3/4) = -.75 right? times x

OpenStudy (anonymous):

just need and answer

OpenStudy (anonymous):

well yeah, but you have to compare (-3/4) to the m of your other equation to see if they are parallel or perpendicular

OpenStudy (anonymous):

I just come here to get quick answers since I have a bunch of these to do and I suck at this.

OpenStudy (anonymous):

they are parallel dude, they both have a slope of -3/4